Jolly Cleaners offers residential and commercial cleaning services. Clients pay a fixed monthly fee for the service, but can cancel the service at the end of any month. In addition to the employees who do the actual cleaning, the firm includes two managers who handle the administrative tasks (human resources, accounting, and so on) and one dispatcher, who assigns the cleaning employees to jobs on a daily basis. On average, residential clients pay $300 per month for cleaning services and the commercial clients pay $1,500 per month. A typical residential client requires 10 hours a month for cleaning and a typical commercial client requires 45 hours a month. In March, Jolly Cleaners had 40 commercial clients and 190 residential clients. Cleaners are paid $12 per hour and are only paid for the hours actually worked. Supplies and other variable costs are estimated to cost $5 per hour of cleaning. Other monthly costs (all fixed) are $51,000 SG&A, including managerial and dispatcher salaries, and $3,100 in other expenses. Jolly Cleaners has earned positive reviews on social media in the area and the managers expect to grow. For April, they forecast a 5 percent increase in residential clients and a 10 percent increase in commercial clients. For July, Jolly Cleaners has budgeted profit of $8,200 based on 60 commercial clients. Required: How many residential clients are budgeted for July? Prepare a budgeted income statement for July to confirm your answer.
Answer #1

Solution:

In July

Number of residential clients = 190 *105% = 199.5

Number of commercial clients = 40 *110 % = 44

BUDGETED INCOME STATEMENT : JULY
Residential Commercial Total
Number of clients:(A) 199.5 44 8,778
Number of hours:(B) 10 45 450
Number of clients - Hours (C) = (A)*(B) 1,995 1,980 3,975
revenue per client (D)($) $300 $1500 $1,800
Total revenue (E) = (D)*(A) $59,850 $66,000 125,850
Cost of cleaner($):(F)= 12*(C) $23,940 $23,760 $47,700
Supplies & other variable cost ($):(G) = 5*(C) $9,975 $9,900 $19,875
Total variable cost($) (H) = (F)+(G) $33,915 $33,660 $67,575
Contribution Margin ($):(J)=(E)-(H) $25,935 $32,340 $58,275
Less: Total fixed costs
SG&A($):(K) $51,000
Other cost ($):(L) $3,100
Total fixed cost($):(M)=(K) +(L) $54,100
Net income ($) = (J) - (M) $4,175